# LLM Reasoning for Cold-Start Item Recommendation

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The application of Large Language Models (LLMs) to recommendation systems has emerged as a promising research direction, drawing significant interest from both academia and industry (carraro2024enhancing; li2024explainable; zhang2023chatgpt). Recent studies highlight the potential of LLM reasoning to further enhance recommendation quality (tsai2024leveraging; liu2025improving). However, these approaches have mainly been evaluated in warm-start scenarios with sufficient historical interactions available for the recommended items.
In practical settings, cold-start scenarios pose a significant challenge for recommender systems, as classic collaborative filtering methods typically underperform due to limited interactions. In contrast, LLMs which equipped from extensive world knowledge and advanced reasoning abilities, are particularly well-positioned to address these challenges by inferring user preferences for new or infrequently interacted items. Despite this potential, current research on LLM for recommendation largely focuses on direct reasoning or reasoning adjusted by supervised fine-tuning, leaving reinforcement learning-based strategies, which are typically considered as more promising and flexible, relatively underexplored.
@@ 2.2.2. Soft Self-Consistency
The second strategy we proposed adapts the Self-Consistency approach (kumar2025llm), which generates multiple parallel reasoning paths and aggregates them for final recommendations. In our implementation, the LLM autonomously constructs diverse reasoning paths based on key aspects of user preferences, deciding which aspects to prioritize. These paths are then synthesized into a final decision. However, rather than employing the classic self-consistency method’s strict majority voting mechanism, we employ a soft summarization strategy, called Soft Self-Consistency, that allows the model to integrate and reconcile its reasoning paths before reaching a conclusion. This modification provides greater autonomy in leveraging reasoning capabilities and can yield performance gains over rigid aggregation policies that risk overlooking nuanced interactions between reasoning perspectives.
@@ 2.3.2. Reinforcement Learning Fine-Tuning
Reinforcement Learning Fine-Tuning (RLFT) has gained significant popularity due to its inherent flexibility in exploring beyond existing training data to achieve superior language generation performance. A critical component of RLFT is the reward function design, which provides supervisory signals to evaluate the quality of LLM-generated responses. In our recommendation context, the reward function should reflect the model’s ability to generate recommendations that align with user preferences. We define our reward function based on parsing the final recommendation output: a reward of 1 is assigned for correct predictions matching the ground truth, -0.1 for incorrect predictions, and -1 for parsing failures. Among the most popular RL methods that are widely adopted in both academia and industry, we select GRPO (shao2024deepseekmath) for its optimal balance between computational efficiency and performance.
@@ 3.4. RL-based Fine-Tuning
Beyond supervised fine-tuning, we conducted experiments to explore the effectiveness of reinforcement learning-based fine-tuning. Specifically, while PPO was the first widely adopted RL method for LLMs with strong performance, it suffers from high computational cost due to its reliance on a large critic model. DPO-based methods simplify training by casting RL as supervised learning but rely on unrealistic assumptions about optimal solutions and have been shown to underperform in practice (xu2024dpo). To address these limitations, we adopt GRPO, a recently proposed RL algorithm that preserves the standard RL framework while improving efficiency by estimating cumulative rewards through sampling, thereby avoiding both the need for a critic model and the shortcomings of DPO.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesisbn: 979-8-4007-2307-0/2026/04doi: 10.1145/3774904.3792872
@@ 1. Introduction
The application of Large Language Models (LLMs) to recommendation systems has emerged as a promising research direction, drawing significant interest from both academia and industry (Carraro and Bridge, 2024; Liu et al., 2024). Recent studies highlight the potential of LLM reasoning to further enhance recommendation quality (Wang et al., 2023; Liu et al., 2025). However, these approaches have mainly been evaluated in warm-start scenarios with sufficient historical interactions available for the recommended items.
In practical settings, cold-start scenarios pose a significant challenge for recommender systems (Li et al., 2021; Yang et al., 2025), as classic collaborative filtering methods typically underperform due to limited interactions. In contrast, LLMs which equipped from extensive world knowledge and advanced reasoning abilities, are particularly well-positioned to address these challenges by inferring user preferences for new or infrequently interacted items. Despite this potential, current research on LLM for recommendation largely focuses on direct reasoning or reasoning adjusted by supervised fine-tuning, leaving reinforcement learning-based strategies, which are typically considered as more promising and flexible, relatively underexplored.
@@ 2.2.2. Soft Self-Consistency
The second strategy we proposed adapts the Self-Consistency approach (Kumar et al., 2025), which generates multiple parallel reasoning paths and aggregates them for final recommendations. In our implementation, the LLM autonomously constructs diverse reasoning paths based on key aspects of user preferences, deciding which aspects to prioritize. These paths are then synthesized into a final decision. However, rather than employing the classic self-consistency method’s strict majority voting mechanism, we employ a soft summarization strategy, called Soft Self-Consistency, that allows the model to integrate and reconcile its reasoning paths before reaching a conclusion. This modification provides greater autonomy in leveraging reasoning capabilities and can yield performance gains over rigid aggregation policies that risk overlooking nuanced interactions between reasoning perspectives.
@@ 2.3.2. Reinforcement Learning Fine-Tuning
Reinforcement Learning Fine-Tuning (RLFT) has gained significant popularity due to its inherent flexibility in exploring beyond existing training data to achieve superior language generation performance. A critical component of RLFT is the reward function design, which provides supervisory signals to evaluate the quality of LLM-generated responses. In our recommendation context, the reward function should reflect the model’s ability to generate recommendations that align with user preferences. We define our reward function based on parsing the final recommendation output: a reward of 1 is assigned for correct predictions matching the ground truth, -0.1 for incorrect predictions, and -1 for parsing failures. Among the most popular RL methods that are widely adopted in both academia and industry, we select GRPO (Shao et al., 2024) for its optimal balance between computational efficiency and performance.
@@ 3.4. RL-based Fine-Tuning
Beyond supervised fine-tuning, we conducted experiments to explore the effectiveness of reinforcement learning-based fine-tuning. Specifically, while PPO was the first widely adopted RL method for LLMs with strong performance, it suffers from high computational cost due to its reliance on a large critic model. DPO-based methods simplify training by casting RL as supervised learning but rely on unrealistic assumptions about optimal solutions and have been shown to underperform in practice (Xu et al., 2024). To address these limitations, we adopt GRPO, a recently proposed RL algorithm that preserves the standard RL framework while improving efficiency by estimating cumulative rewards through sampling, thereby avoiding both the need for a critic model and the shortcomings of DPO.